OpenStudy (anonymous):

Find the perimeter of a rectangle Base= (x+2)in Height=(x-1)in

OpenStudy (anonymous):

use 2(b+h)=2((x+2)+(x-1))=2(2x+1)=4x+2

OpenStudy (anonymous):

what is x?